OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: 9x2 − 6x + 1 36b2 − 24b + 8 16x2 + 24x + 9 4a2 − 10a + 25 PLEASE HELP!:)

OpenStudy (compassionate):

No one snipe me.

OpenStudy (anonymous):

lol

OpenStudy (compassionate):

Here are some examples: a) 2x² + 7x − 15 = (2x − 3)(x + 5) b) 2x² − 7x − 15 = (2x + 3)(x − 5) c) 2x² − x − 15 = (2x + 5)(x − 3) d) 2x² − 13x + 15 = (2x − 3)(x − 5)

OpenStudy (anonymous):

But how do you tell if they are perfect square trinomials?

OpenStudy (compassionate):

The square of a binomial (a + b)² <=== Form The square of a binomial comes up so often that the student should be able to write the final product immediately. It will turn out to be a very specific trinomial. To see that, let us square (a + b): (a + b)² = (a + b)(a + b) = a² + 2ab + b². For, the outers plus the inners will be ab + ba = 2ab. The order of factors does not matter. The square of any binomial produces the following three terms: 1. The square of the first term of the binomial: a² 2. Twice the product of the two terms: 2ab 3. The square of the second term: b² The square of every binomial has that form: a² + 2ab + b². To recognize that is to know an essential product in the "multiplication table" of algebra. Is this a perfect square trinomial: x² + 14x + 49 ? Answer: Yes. It is the square of (x + 7). x² is the square of x. 49 is the square of 7. And 14x is twice the product of x· 7. In other words, x² + 14x + 49 could be factored as x² + 14x + 49 = (x + 7)²

OpenStudy (anonymous):

ok so that means b, c, and d are correct?

OpenStudy (compassionate):

Yes, sir! See, it's easy; isn't it? ==============================> BEST ANSWER?

OpenStudy (anonymous):

lol thanks

OpenStudy (compassionate):

No problem.